# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 4.1 Preliminary results for the analysis
Because $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, and $\tilde{\psi}_{p_{k+1}}(x^{k+1},y^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k+1},y^{k+1})$ following from the non-decreasing property of the penalty parameter $p_{k}$ and the non-negative property of $F-\underline{F}$, we can obtain the conclusion from (37). ∎
@@ 4.2 Convergence to stationary points
In this part, we establish the sequential convergence properties of the AGILS, described in 1. Throughout this subsection, we assume that Assumptions 1, 2, and 3 hold, and $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$. The sequences $\{(x^{k},y^{k})\}$ and $\{p_{k}\}$ are generated by the algorithm, where the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants, and $s_{k}$ satisfies $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$. Our analysis demonstrates the sequential convergence of $\{(x^{k},y^{k})\}$ to a stationary point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
We first show that the differences between consecutive iterates are square summable.
Lemma 6.
The sequence $\left\{(x^{k},y^{k})\right\}$ satisfies
$\sum\limits_{k=1}^{{\infty}}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|^{2}<\infty,\quad\text{and}\quad\lim\limits_{k\to\infty}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|=0.$
@@
Summing 26 from Lemma 5 over $k=0,1,\dots,K-1$ yields,
$\tilde{\psi}_{p_{K}}\left(x^{K},y^{K}\right)-\tilde{\psi}_{p_{0}}\left(x^{0},y^{0}\right)\leq$ $-\sum\limits_{k=0}^{K-1}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}\right)$ (38)
$+\sum\limits_{k=0}^{K-1}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}.$
Since $\{s_{k}\}$ is square summable, it follows that
$\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}<\infty.$
Additionally, because $\tilde{\psi}_{p_{K}}\left(x^{K},y^{K}\right)\geq 0$, by taking $K\rightarrow\infty$ in (38), we have
$\quad\,\sum\limits_{k=0}^{\infty}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-y^{k}\right\|^{2}\right)$ (39)
$\leq\tilde{\psi}_{p_{0}}\left(x^{0},y^{0}\right)+\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}s_{k}^{2}<\infty,$
which proves the result. ∎
We now present the main result, showing that every accumulation point of sequence ${(x^{k},y^{k})}$ is a KKT point of the problem $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
Theorem 2.
Let $\epsilon>0$, if the sequence $\{p_{k}\}$ is bounded, then every accumulation point $(\bar{x},\bar{y})$ of $\{(x^{k},y^{k})\}$ is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
Proof.
The proof builds upon techniques used in the proof of Theorem 1 in [27]. Let $(\bar{x},\bar{y})$ be an accumulation point of the sequence $\{(x^{k},y^{k})\}$, and let $\{(x^{k_{j}},y^{k_{j}})\}$ be the corresponding subsequence such that $(x^{k_{j}},y^{k_{j}})\rightarrow(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. Since $\{p_{k}\}$ is bounded, there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$. From 6, we know that
$\lim\limits_{k\to\infty}\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|=0,$ (40)
which implies that for sufficiently large $k$, the inequality $\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|\leq c_{p}/p_{k}$ holds. According to the update strategy of $p_{k}$ given in (23), we have, for large $k$, that $t^{k}<c_{p}^{-1}\|(x^{k},y^{k})-(x^{k-1},y^{k-1})\|$ and thus $t^{k}\rightarrow 0$. Consequently, from the definition of $t^{k}$, we obtain
$\max\left\{\varphi(x^{k},y^{k})-\varphi(x^{k},\theta^{k})-\frac{1}{2\gamma}\|\theta^{k}-y^{k}\|^{2}-\epsilon,0)\right\}\rightarrow 0.$ (41)
Next, using the update rules for $x^{k+1}$ and $y^{k+1}$ given in (20) and (15), and the expressions of $d_{x}^{k}$ and $d^{k}_{y}$, we have
@@
$\xi_{y}^{k}$ $:=p_{k}\left(\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\right)-\frac{p_{k}}{\beta_{k}}\left(y^{k+1}-y^{k}\right).$

### v2

@@ 4.1 Preliminary results for the analysis
Because $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, and $\tilde{\psi}_{p_{k+1}}(x^{k+1},\tilde{y}^{k+1})\leq\tilde{\psi}_{p_{k}}(x^{k+1},\tilde{y}^{k+1})$ following from the non-decreasing property of the penalty parameter $p_{k}$ and the non-negative property of $F-\underline{F}$, we can obtain (30) from (41).
Finally, we show that $\{\zeta_{k}\}$ is square summable. Let $\mathcal{G}_{0}=\mathcal{G}(\theta^{0},x^{0},y^{0})$ and $r_{k}:=(\prod_{i=0}^{k}\tau_{i})\mathcal{G}_{0}$. Since $\zeta_{k}:=\max\{s_{k},r_{k}\}$, we have $\sum_{k=0}^{\infty}\zeta_{k}^{2}\leq\sum_{k=0}^{\infty}s_{k}^{2}+\sum_{k=0}^{\infty}r_{k}^{2}$. By assumption, $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$. Thus, we only need to show that $\{r_{k}\}$ is square summable. Recall that $\tau_{k}=\tau_{0}/(k+1)^{p_{\tau}}$ with $p_{\tau}>0$, then
$r_{k}=\left(\prod_{i=0}^{k}\tau_{i}\right)\mathcal{G}_{0}=\tau_{0}^{k+1}\left(\prod_{i=1}^{k+1}\frac{1}{i^{p_{\tau}}}\right)\cdot\mathcal{G}_{0}=\frac{\tau_{0}^{k+1}}{((k+1)!)^{p_{\tau}}}\mathcal{G}_{0}.$
For the ratio test on $\sum_{k=0}^{\infty}r_{k}^{2}$, we consider
$\lim_{k\rightarrow\infty}\frac{r_{k+1}^{2}}{r_{k}^{2}}=\lim_{k\rightarrow\infty}\frac{\tau_{0}^{2}}{(k+2)^{2p_{\tau}}}=0.$
Therefore, by the ratio test, $\sum_{k=0}^{\infty}r_{k}^{2}<\infty$, implying that $\{\zeta_{k}\}$ is square summable. ∎
4.2 Analysis of the feasibility correction
Throughout this subsection and the remainder of this section, we assume that Assumptions 1, 2, and 3 hold, and $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$. The sequences $\{(x^{k},y^{k},\tilde{y}^{k})\}$ and $\{p_{k}\}$ are generated by AGILS (Algorithm 1), where the step sizes satisfy $\alpha_{k}\in[\underline{\alpha},2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[\underline{\beta},2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $\underline{\alpha},\underline{\beta},c_{\alpha},c_{\beta}$ are positive constants. The inexactness parameters satisfy $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$ and $\tau_{k}:=\tau_{0}/(k+1)^{p_{\tau}}$ with $\tau_{0}>0$ and $p_{\tau}>0$.
In this subsection, we show that the feasibility correction procedure (Algorithm 2) enforces the feasibility of the generated iterates. We begin by showing that the differences between consecutive iterates are square summable.
Lemma 6.
The sequence $\left\{(x^{k},y^{k},\tilde{y}^{k})\right\}$ satisfies
$\sum\limits_{k=0}^{{\infty}}\left\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\right\|^{2}<\infty,\quad\text{and}\quad\lim\limits_{k\to\infty}\left\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\right\|=0.$
Proof.
Summing (30) from Lemma 5 over $k=0,1,\dots,K-1$ yields,
$\tilde{\psi}_{p_{K}}\left(x^{K},\tilde{y}^{K}\right)-\tilde{\psi}_{p_{0}}\left(x^{0},\tilde{y}^{0}\right)\leq$ $-\sum\limits_{k=0}^{K-1}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}\right)$ (42)
$+\sum\limits_{k=0}^{K-1}\left(\frac{C_{\eta}^{2}}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{C_{\eta}^{2}}{c_{\beta}\gamma^{2}}\right)\zeta_{k}^{2}+\sum\limits_{k=0}^{K-1}C_{\gamma}\zeta_{k+1}^{2}.$
Since $\{\zeta_{k}\}$ is square summable from Lemma 5, it follows that
$\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}+\sum\limits_{k=0}^{\infty}C_{\gamma}\zeta_{k+1}^{2}<\infty.$
Additionally, because $\tilde{\psi}_{p_{K}}\left(x^{K},\tilde{y}^{K}\right)\geq 0$, by taking $K\rightarrow\infty$ in (42), we have
$\quad\,\sum\limits_{k=0}^{\infty}\left(\frac{c_{\alpha}}{4}\left\|x^{k+1}-x^{k}\right\|^{2}+\frac{c_{\beta}}{4}\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}\right)$ (43)
$\leq\tilde{\psi}_{p_{0}}\left(x^{0},\tilde{y}^{0}\right)+\sum\limits_{k=0}^{\infty}\left(\frac{1}{c_{\alpha}}(L_{f_{x}}+L_{g_{2}})^{2}+\frac{1}{c_{\beta}\gamma^{2}}\right)C_{\eta}^{2}\zeta_{k}^{2}+\sum\limits_{k=0}^{\infty}C_{\gamma}\zeta_{k+1}^{2}<\infty,$
which proves the result. ∎
The following lemma shows that any accumulation point of the corrected iterates $\tilde{y}^{k}$ satisfying condition (24) lies in the solution set of the lower-level problem.
Lemma 7.
Let $\{(x^{k_{j}},\tilde{y}^{k_{j}})\}$ be any convergent subsequence of iterates $\{(x^{k},\tilde{y}^{k})\}$ such that $(x^{k_{j}},\tilde{y}^{k_{j}})$ satisfies condition (24) with $k=k_{j}-1$. If $(x^{k_{j}},\tilde{y}^{k_{j}})\rightarrow(\bar{x},\bar{y})$ as $j\rightarrow\infty$ for some $(\bar{x},\bar{y})$, then $\bar{y}\in S(\bar{x})$.
Proof.
Define the residual term $\tilde{r}^{k}:=\tilde{y}^{k}-\operatorname{Prox}_{\tilde{g}(x^{k},\cdot)}\left(\tilde{y}^{k}-\nabla_{y}f(x^{k},\tilde{y}^{k})\right)$. By the first-order optimality condition for the proximal operator, $\tilde{r}^{k}$ satisfies
$\tilde{r}^{k}\in\nabla_{y}f(x^{k},\tilde{y}^{k})+\partial_{y}\tilde{g}(x^{k},\tilde{y}^{k}-\tilde{r}^{k}).$ (44)
The sequence $(x^{k_{j}},\tilde{y}^{k_{j}})$ satisfies condition (24) with $k=k_{j}-1$, which implies
$\left\|\tilde{r}^{k_{j}}\right\|\leq\frac{c_{\tilde{y}}}{p_{k_{j}}}\left\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\right\|.$
From Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. As $p_{k_{j}}\geq p_{0}$, the above inequality ensures that $\lim_{k_{j}\to\infty}\|\tilde{r}^{k_{j}}\|=0$.
Taking the limit as $k_{j}\rightarrow\infty$ in (44) with $k=k_{j}$, by the continuity of $\nabla_{y}f$ and the outer semicontinuity of $\partial_{y}\tilde{g}$, it follows that
$0\in\nabla_{y}f(\bar{x},\bar{y})+\partial_{y}\tilde{g}(\bar{x},\bar{y}).$
Since the lower-level problem is convex in $y$, this condition implies $\bar{y}\in S(\bar{x})$. ∎
We next establish the feasibility of the iterates generated by AGILS with respect to the constraint $\varphi({x},{y})-v_{\gamma}({x},{y})\leq\epsilon$ when the penalty parameter $p_{k}$ is bounded.
Proposition 3.
Let $\epsilon>0$. If the sequence $\{p_{k}\}$ is bounded, then any accumulation point $(\bar{x},\bar{y})$ of the iterate sequence $\{(x^{k},{y}^{k})\}$ satisfies
$\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon.$
Proof.
Let $(\bar{x},\bar{y})$ be an accumulation point of $\{(x^{k},{y}^{k})\}$, and let $\{k_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N}$ be an index subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ as $j\rightarrow\infty$. By Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})$ also converges to $(\bar{x},\bar{y})$
Since the sequence $\{p_{k}\}$ is bounded and non-decreasing by construction in AGILS (Algorithm 1), there exists $k_{0}>0$ such that $p_{k}=p_{k_{0}}$ for all $k\geq k_{0}$. By passing to a subsequence if necessary, we can assume that $k_{j}-1\geq k_{0}$ for all $j\in\mathbb{N}$.
According to the update rule of $p_{k}$ in AGILS (Algorithm 1), the condition $p_{k_{j}-1}=p_{k_{j}}$ implies that one of the following two cases must occur for each $k=k_{j}-1$:
Case I: $\left\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\right\|\geq c_{p}\min\left\{1/p_{k_{0}},{t^{k_{j}}}\right\}$ and $\tilde{y}^{k_{j}}=y^{k_{j}}$.
Case II: $\left\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\right\|<c_{p}\min\left\{1/p_{k_{0}},{t^{k_{j}}}\right\}$ and the feasibility correction procedure is applied to generate $\tilde{y}^{k_{j}}$ satisfying condition (24).
At least one of these two cases must hold for an infinite number of indices in the subsequence $\{k_{j}-1\}$. We analyze them separately.
Case 1: Suppose Case I occurs for an infinite subsequence of indices. By passing to a further subsequence, we assume it holds for all $j\in\mathbb{N}$.
The condition $\tilde{y}^{k_{j}}=y^{k_{j}}$ combined with $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ implies that $(x^{k_{j}},{y}^{k_{j}})\to(\bar{x},\bar{y})$. From Lemma 6, we know $\lim_{k_{j}\to\infty}\|(x^{k_{j}},y^{k_{j}})-(x^{k_{j}-1},\tilde{y}^{k_{j}-1})\|=0$. For the condition of Case I to hold, we must therefore have $t^{k_{j}}\rightarrow 0$. This implies
$\limsup_{j\rightarrow\infty}\left\{\varphi(x^{k_{j}},y^{k_{j}})-\varphi(x^{k_{j}},\theta^{k_{j}})-\frac{1}{2\gamma}\left\|\theta^{k_{j}}-y^{k_{j}}\right\|^{2}-\epsilon\right\}\leq 0.$
Since $\zeta_{k}\rightarrow 0$ and $\mathcal{G}(\theta^{k_{j}},x^{k_{j}},y^{k_{j}})\leq\zeta_{k_{j}}$, Lemma 2 ensures that $\|\theta^{k_{j}}-\theta^{*}_{\gamma}(x^{k_{j}},y^{k_{j}})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2), it follows that $\theta^{k_{j}}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}},\theta^{k_{j}})+\frac{1}{2\gamma}\|\theta^{k_{j}}-y^{k_{j}}\|^{2}\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Then taking the limit in the inequality above as $j\rightarrow\infty$ yields $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon\leq 0$.
Case 2: Now, suppose Case II occurs for an infinite subsequence of indices. Again, we assume it holds for all $j\in\mathbb{N}$.
In this scenario, the feasibility correction procedure is applied at each iteration $k=k_{j}-1$ to generate $\tilde{y}^{k_{j}}$ satisfying condition (24). Since $(x^{k_{j}},\tilde{y}^{k_{j}})\rightarrow(\bar{x},\bar{y})$ and satisfies condition (24), it follows from Lemma 7 that $\bar{y}\in S(\bar{x})$. For any such point, $\theta^{*}_{\gamma}(\bar{x},\bar{y})=\bar{y}$ (see, e.g., [7, Theorem 10.7]), and therefore $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})=0$. Since $\epsilon>0$, the constraint $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$ is satisfied.
Since any accumulation point must be the limit of a subsequence that falls into one of these two cases, we conclude that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\leq\epsilon$. ∎
The following theorem establishes that the penalty parameter sequence $\left\{p_{k}\right\}$ is bounded, provided the sequence of iterates $\left\{(x^{k},y^{k})\right\}$ is bounded.
Theorem 1.
Let $\epsilon>0$. If the sequence of iterates $\left\{(x^{k},y^{k})\right\}$ is bounded, then the penalty parameter sequence $\left\{p_{k}\right\}$ must be bounded.
Proof.
The proof is by contradiction. Assume that $p_{k}\to\infty$ as $k\to\infty$. According to the penalty update rule in AGILS (Algorithm 1), this implies the penalty parameter is increased at infinitely many iterations. Let $\{k_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N}$ be the subsequence of iteration indices where such an update occurs.
Since $\{(x^{k},y^{k})\}$ is bounded and Lemma 6 establishes that $\lim_{k\to\infty}\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\|=0$, the sequence $\{\tilde{y}^{k}\}$ is also bounded. By passing to a subsequence if necessary, we can assume that for the indices $\{k_{j}\}$, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ for some $(\bar{x},\bar{y})\in X\times Y$ and $\tilde{y}^{k_{j}+1}\to\tilde{y}$ for some $\tilde{y}\in Y$. From Lemma 6, it also follows that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$
At each iteration $k=k_{j}$, the penalty parameter is increased, which requires that condition (22) is met
$t^{k_{j}+1}>c_{p}^{-1}\left\|\left(x^{k_{j}+1},{y}^{k_{j}+1}\right)-\left(x^{k_{j}},\tilde{y}^{k_{j}}\right)\right\|>0.$ (45)
We will show that this leads to the conclusion that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$. Let us assume the contrary, i.e., $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$. From Lemma 2, the conditions $\mathcal{G}(\theta^{k_{j}+1},x^{k_{j}+1},y^{k_{j}+1})\leq\zeta_{k_{j}+1}$ and $\zeta_{k}\rightarrow 0$ ensure that $\|\theta^{k_{j}+1}-\theta^{*}_{\gamma}(x^{k_{j}+1},y^{k_{j}+1})\|\rightarrow 0$. By the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2) and $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$, we have $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. The continuity of $\varphi$ then implies that $\varphi(x^{k_{j}+1},\theta^{k_{j}+1})+\|\theta^{k_{j}+1}-y^{k_{j}+1}\|^{2}/(2\gamma)\rightarrow v_{\gamma}(\bar{x},\bar{y})$. Our assumption $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})-\epsilon<0$, combined with the continuity of $\varphi$, means that for $j$ sufficiently large,
$\varphi(x^{k_{j}+1},y^{k_{j}+1})-\varphi(x^{k_{j}+1},\theta^{k_{j}+1})-\frac{1}{2\gamma}\left\|\theta^{k_{j}+1}-y^{k_{j}+1}\right\|^{2}-\epsilon<0.$
This inequality implies that $t^{k_{j}+1}=0$ which contradicts (45). Thus, our assumption was false, and we must have $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$.
An increase in the penalty parameter at iteration $k=k_{j}$ corresponds to one of the following two cases:
Case I: $\left\|y^{k+1}-\theta^{k+1}\right\|<c_{y}\gamma/p_{k}$.
Case II: $\left\|y^{k+1}-\theta^{k+1}\right\|\geq c_{y}\gamma/p_{k}$, the feasibility correction procedure is applied, but the corrected iterate $\tilde{y}^{k+1}$ violates the descent condition (25).
At least one of these cases must occur for an infinite subsequence of $\{k_{j}\}$. We analyze both possibilities to derive a contradiction.
Case 1: Assume an infinite subsequence, still denoted by $\{k_{j}\}$, falls into Case I. For this subsequence, we have
$\left\|y^{k_{j}+1}-\theta^{k_{j}+1}\right\|<\frac{\gamma}{p_{k_{j}+1}}\to 0.$
Since we assume $p_{k_{j}}\rightarrow\infty$, the right-hand side converges to 00. As $j\to\infty$, we have $y^{k_{j}+1}\rightarrow\bar{y}$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Taking the limit yields $\|\bar{y}-\theta^{*}_{\gamma}(\bar{x},\bar{y})\|\leq 0$, which implies $\bar{y}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. By definition, this yields $\varphi(\bar{x},\bar{y})=v_{\gamma}(\bar{x},\bar{y})$, which contradicts our established result that $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$.
Case 2: Assume an infinite subsequence, again denoted by $\{k_{j}\}$, belongs to Case II. Here, the feasibility correction procedure is applied at each iteration $k=k_{j}$, generating a corrected iterate $\tilde{y}^{k_{j}+1}$ satisfying condition (24) and a corresponding $\tilde{\theta}^{k_{j}+1}$ satisfying the inexact criterion (26). Consequently, we have $\mathcal{G}(\tilde{\theta}^{k_{j}+1},x^{k_{j}+1},\tilde{y}^{k_{j}+1})\leq\zeta_{k_{j}+1}$. However, these iterates violate the descent condition (25), leading to
$\frac{1}{p_{k_{j}}}F(x^{k_{j}+1},\tilde{y}^{k_{j}+1})+\varphi(x^{k_{j}+1},\tilde{y}^{k_{j}+1})-\varphi(x^{k_{j}+1},\tilde{\theta}^{k_{j}+1})-\frac{1}{2\gamma}\|\tilde{\theta}^{k_{j}+1}-\tilde{y}^{k_{j}+1}\|^{2}$
$>$ $\frac{1}{p_{k_{j}}}F(x^{k_{j}+1},y^{k_{j}+1})+\varphi(x^{k_{j}+1},y^{k_{j}+1})-\varphi(x^{k_{j}+1},\theta^{k_{j}+1})-\frac{1}{2\gamma}\|\theta^{k_{j}+1}-y^{k_{j}+1}\|^{2}.$
As $j\rightarrow\infty$, we have $(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\rightarrow(\bar{x},\tilde{y})$. Since $(x^{k_{j}+1},\tilde{y}^{k_{j}+1})$ satisfies condition (24), it follows from Lemma 7 that $\tilde{y}\in S(\bar{x})$. Moreover, given that $\zeta_{k}\rightarrow 0$ and $\mathcal{G}(\tilde{\theta}^{k_{j}+1},x^{k_{j}+1},\tilde{y}^{k_{j}+1})\leq\zeta_{k_{j}+1}$, Lemma 2 ensures that $\|\tilde{\theta}^{k_{j}+1}-\theta^{*}_{\gamma}(x^{k_{j}+1},\tilde{y}^{k_{j}+1})\|\rightarrow 0$, and thus $\tilde{\theta}^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\tilde{y})$ from the continuity of $\theta^{*}_{\gamma}(x,y)$ (Proposition 2).
Because $\tilde{y}\in S(\bar{x})$, we know that $\theta^{*}_{\gamma}(\bar{x},\tilde{y})=\tilde{y}$ (see, e.g., [7, Theorem 10.7]). Taking the limit of the left-hand side of the inequality above as $k_{j}\rightarrow\infty$, and using the continuity of $F$ and $\varphi$ along with $p_{k_{j}}\rightarrow\infty$, we find that the expression converges to $\varphi(\bar{x},\tilde{y})-\varphi(\bar{x},\theta^{*}_{\gamma}(\bar{x},\tilde{y}))-\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(\bar{x},\tilde{y})-\tilde{y}\|^{2}=0$. On the other hand, we have $(x^{k_{j}+1},y^{k_{j}+1})\to(\bar{x},\bar{y})$ and $\theta^{k_{j}+1}\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, the right-hand side of the inequality converges to $\varphi(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon>0$. As $k_{j}\rightarrow\infty$, the inequality thus implies $00\geq\epsilon>$ 0, which is a contradiction.
Since both possible scenarios for an infinite number of penalty updates lead to a contradiction, our initial assumption that $p_{k}\to\infty$ must be false. Therefore, the sequence of penalty parameters $\{p_{k}\}$ is bounded.
@@ 4.3 Convergence to stationary points
In this part, we establish the sequential convergence properties of the AGILS (Algorithm 1). We demonstrates that any accumulation point of sequence $\{(x^{k},y^{k})\}$ is a KKT point of the problem $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$. The problem setting and algorithmic assumptions are identical to those in the previous subsection.
Theorem 2.
Let $\epsilon>0$. If the sequence $\{p_{k}\}$ is bounded, then any accumulation point $(\bar{x},\bar{y})$ of $\{(x^{k},y^{k})\}$ is a KKT point of $({\rm VP})_{\gamma}^{\bar{\epsilon}}$ for some $\bar{\epsilon}\leq\epsilon$.
@@
The proof builds upon techniques used in the proof of Theorem 1 in [61]. Let $(\bar{x},\bar{y})$ be an accumulation point of the sequence $\{(x^{k},y^{k})\}$, and let $\{(x^{k_{j}+1},y^{k_{j}+1})\}$ be a convergent subsequence such that $(x^{k_{j}+1},y^{k_{j}+1})\rightarrow(\bar{x},\bar{y})$ as $k_{j}\rightarrow\infty$. By Lemma 6, we have $\lim_{k_{j}\to\infty}\|(x^{k_{j}+1},y^{k_{j}+1})-(x^{k_{j}},\tilde{y}^{k_{j}})\|=0$, which implies that $(x^{k_{j}},\tilde{y}^{k_{j}})\to(\bar{x},\bar{y})$ Since $\{p_{k}\}$ is bounded, there exists $k_{0}>0$ such that $p_{k}=\bar{p}:=p_{k_{0}}$ for any $k\geq k_{0}$.
From the update rules (19) and (14), and the expressions for $d_{x}^{k}$ and $d_{y}^{k}$, we have
@@
$\xi_{y}^{k}$ $:=p_{k}\left(\nabla_{y}\left(\tilde{\psi}_{p_{k}}-g\right)\left(x^{k},y^{k+1}\right)-d_{y}^{k}\right)-\frac{p_{k}}{\beta_{k}}\left(y^{k+1}-\tilde{y}^{k}\right).$